Determine the forces in members AB, AC and (additionally) CD? State whether it is compression or tension. THANK YOU!!!
Determine the forces in members AB and AC. 150 lb 100 lb Αό- D B 7 3 C brod 4 4-
Answer #1

Complete solution is given below.

solution : ZMD = 0 - (150x$) - (100 x 4) + (Re * 4) - 0 (Re = 400 lb E Fy=0 150 + 100 Rct by (Oy -150 lb E Fx = 0 Dx = 0 cons* Consider Joint D - FOR Foo och -150 FCD EFY =0 -150 Fcp sino=0 (Fco 250 lb compression